# The trimer paradox: the effect of stiff constraints on equilibrium distributions in overdamped dynamics

### v2

@@ I Introduction
In many applications, the dynamics of the hard degrees of freedom are of secondary importance, and one method of circumventing this difficulty is to treat them as fully constrained (i.e. rigid), thus eliminating the short timescale, and enabling faster simulations. This is the basis of many algorithms, such as SHAKE [4] and its further extensions. During the development of such rigid models, a disconnect was discovered between the equilibrium bond angle distributions in the rigid simulations and the simulations with bonds modelled as very stiff springs.
@@
In this contribution, we present a clear mathematical procedure leading to the aforementioned limits in the overdamped regime when inertial effects can be neglected. We show that the distribution given by Eq. (1) is not completely determined by fixing the bond lengths. To correctly predict the limiting distribution, knowledge about the nature of confinement is needed, and its structure survives in the final expression. Furthermore, we demonstrate that Fixman’s expressions for the stiff limit miss a term, which is important even in the case of harmonic springs.
@@
Fixman [5] presented a purely classical calculation of the limiting distribution, correctly writing the action of the distribution on a test function but erroneously assuming that the constraining potential depends only on the hard coordinates. As a result, the details of the confining potential cannot appear in his final expressions, which contain only the projected phase-space volume element terms. Gō and Sheraga [6] treated the problem quantum-mechanically and correctly identified two contributions to the equilibrium probability density of the soft coordinates: the projected volume elements and the zero-point energy of the vibrational motion of the hard coordinates. However, they then stated that the zero-energy contribution in typical molecular systems varies only a little when changing the soft coordinates and may be altogether neglected. We show in the following that this is not necessarily the case in overdamped dynamics. By focusing on a concrete classical two-dimensional example with one constraint, Helfand [11] correctly identified the rigid-rod type distribution as a uniform distribution on the constraining manifold and computed the stiff-spring limit in a similar simple case. Unfortunately, these results are not expressed in a form applicable to higher-dimensional examples with multiple constraints.
@@
Hinch et al [8, 12] approached the problem differently, by manipulating the Langevin equations, which are notoriously difficult to handle, and erroneously assumed that details of the confining potential are of little importance to the equilibrium distribution of the constrained configurations. Van Kampen and Lodder [13] noted that the approach of Helfand [11] is generally applicable to constrained systems and comment that the discrepancy between the stiff-spring and rigid-rod distributions in the case of the trimer molecule is due to the fact that the width of this gully is not the same everywhere [13]. However, they did not quantify the effect of this width in a higher-dimensional case. As a result, the knowledge of the influence of the shape of the constraining potential appears to have been lost as multiple works (for example, [14, 15, 3]) mention only the densities of the volume elements and omit the zero point energy or, equivalently, the potential shape terms from their descriptions of constrained dynamics, widely using the result of Fixman [5] as a reference.
In the following, we show a way to compute the limiting distribution by integrating out the hard degrees of freedom in the overdamped limit. We explicitly find the metric contribution arising from the transformation of coordinates and the shape term that represents the details of the constraining potential and which remains imprinted on the configurational distribution when the hard coordinates are taken to be infinitely stiff.
@@ III The stiff spring limit
This work concerns two distributions – the rigid rod distribution and the distributional limit of equilibrium distributions at increasingly stiff confinement: the stiff spring limit. We begin our consideration with a careful definition of the stiff spring limit with the rigid rod distribution discussed in later sections.
We present a derivation of the limiting distribution in the following setting, sketched in Fig. 1. Consider a molecule (or a polymer) with a handful of identical subunits (atoms, beads, or monomers) each of mass $m$ and a total of $2N$ degrees of freedom ($N$ configurational and $N$ momentum) in a heat bath at constant temperature $T$, with $\beta=1/k_{B}T$. The conformation of the molecule is described by an $N$-dimensional vector $\bm{q}$ and the velocities of monomers are specified by a vector $\bm{v}$. We parametrise the conformations by Cartesian coordinates of monomers $q_{i}$ and Cartesian components of velocities $v_{i}$ with $i\in 1\dots N$. The potential energy of the molecule has two components: the conformation-dependent energy $U(\bm{q})$ and the confining (springs) potential $k^{2}W(\bm{q})$ where $k$ is a large parameter describing the spring stiffness. Note that $k$ here would correspond to the square root of a harmonic spring constant. The probability density $\hat{p}_{k}$ of the equilibrium distribution of the system configuration with respect to the Lebesgue measure $\mathrm{d}\bm{q}\mathrm{d}\bm{v}$ is given by the Maxwell-Boltzmann distribution [16]
@@
with a (different) normalising function $N(k)$. Crucially, this density is given with respect to canonical conformation coordinates and not just any coordinates – if generalised coordinates are used, more care is needed and volume element density terms appear in expressions analogous to Eq. (4). Moreover, in the case of generalised coordinates, the expression for kinetic energy is also more complicated. Clearly, Eq. (4) is a physical object and we are free to work in any coordinates of choice, and for the purpose of the derivation of the stiff spring limit Cartesian coordinates are particularly convenient. We return to the treatment of generalized coordinates in later chapters of this manuscript.
@@
$I_{k}=\iint\frac{N(k)}{k}\exp\left\{-\beta U(\bm{r},k^{-1}\bm{t})-\beta\left[\bm{t}^{T}\bm{\mathsf{H}}(\bm{r})\bm{t}+||\bm{t}||^{2}\frac{o(k^{-2}||\bm{t}||^{2})}{k^{-2}||\bm{t}||^{2}}\right]\right\}\phi(\bm{r},k^{-1}\bm{t})\mathrm{d}\bm{t}\mathrm{d}\bm{r}.$ (9)
@@
$I=\iint\left(\lim_{k\to\infty}\frac{N(k)}{k}\right)\exp\left[-\beta U(\bm{r},\bm{0})-\beta\,\bm{t}^{T}\bm{\mathsf{H}}(\bm{r})\bm{t}\right]\phi(\bm{r},\bm{0})\mathrm{d}\bm{t}\mathrm{d}\bm{r},$ (10)
@@
$L=\lim_{k\to\infty}\frac{N(k)\sqrt{(2\pi\beta)^{N-M}}}{k},$ (12)
@@
for an appropriate constant $L$, and since $\phi$ is arbitrary we arrive at $p_{\infty}=\lim_{k\to\infty}p_{k}$ such that
@@
in the weak sense. We have thus arrived at the central result of this manuscript.
@@
III.1 Calculation of the shape term
Figure 2: (a) Parametrisation of the trimer uses the position of the central particle ($b$), bond angle $\psi$ and bond lengths $l_{a}$ and $l_{b}$, with identical equilibrium spring lengths $l_{0}$. Rigid-body rotation around the particle $b$ is omitted for clarity. (b) Motion along the symmetric (top) and antisymmetric eigenvectors of the hessian $\bm{\mathsf{H}}$ of the harmonic confinement function. Note that the central bead must move for the overall motion to be orthogonal to the constraining manifold $\mathcal{M}$. (c) Equilibrium probability density function (PDF) for the bond angle $\psi$ from the Langevin numerical simulation. Solid line is the theoretical prediction of Eq. (34).
@@ III.1 Calculation of the shape term
IV Calculation for the trimer
@@
from the matrix $\bm{\mathsf{J}}$ listed in the Appendix. Next, we compute the shape term using
@@
where $N$ is a normalisation constant. Therefore, the marginal density of the bond angle is proportional to $\sin\psi$ as expected. However, this is simply not the case of a uniform distribution––that distribution corresponds to $|\bm{\mathsf{H}}|=\text{const}$ and the notion of uniformity is inherited from the canonical measure in the canonical (Cartesian) coordinates. This can be examined in greater detail by looking at the eigenvectors of $\bm{\mathsf{H}}$, visualised in Fig. 2(b). There are two eigenvectors which correspond to symmetric and antisymmetric motion of the terminal beads. Counterintuitively, the vectors in $\mathcal{M}^{\perp}$ change the orientation of at least one bond when the bond length changes.
Figure 3: (a) The shape of a cyclic tetramer is specified with the bead $b$ taken as reference point. The directions of sections $ab$ and $bc$ span a rhombus, which is used to define the position of bead $d$. (b) Equilibrium probability density function (PDF) for the bond angle $\psi$ from the numerical simulation. Solid line is the theoretical prediction of Eq. (49).
@@
V Cyclic tetramer
@@
VI Conclusions
In this work, we have provided a procedure by which one can find the equilibrium probability probability distribution function of the configurations of a bead-spring model of a polymer in the overdamped regime in the limit of infinite stiffness of the bonds. In particular, we have demonstrated that the shape of the confining spring potential persists in the limiting distribution. By rephrasing the problem as a mathematically rigorous limit, we have shown that the classical expression given by Fixman [5], later reproduced in numerous works and books, for example in Ref. [3], is missing a critical term describing the shape of the confining potential, $|\bm{\mathsf{H}}|$.
This shows that the reasoning presented by Frenkel & Smit [3] cannot be universally applied. First, the dynamics and the distribution of a polymer connected with (stiff or soft) springs is independent of the mass of the monomers in the overdamped regime, and thus the limiting distribution cannot depend on the mass of the monomers. Second, putting aside the question of existence of the rigid rod distribution (or the lack thereof), the supposition that the stiff spring limit should always be the same, regardless of the details of the confining potential is clearly not true, and harmonic potentials will lead to a different outcome than springs that realise uniform confinement around the constraining manifold, as previously remarked, e.g., by van Kampen & Lodder [13], although without a general mathematical formulation. The difference can be arbitrarily large, as shown in the tetramer example. Frenkel & Smit remark in their book that for the bond length constraints of the type most often used in MD simulations, the effect of hard constraints on the distribution functions seems to be relatively small [3, Chapter 15.1.1]. We believe this statement to be misleading. In the discussed case of the trimer molecule, the effect might indeed be small, but we have shown that for the tetramer system, which is still a relatively simple setup, the effect is pronounced. To provide a practical insight, we propose a straightforward technique for the analysis of the limiting distributions once the confining potential is specified in detail.
@@ Acknowledgements.
The Authors would like to thank Bartłomiej Lewandowski for his insightful comments regarding convergence of distributions. The work of ML and RW was supported by the National Science Centre of Poland (FundRef DOI: http://dx.doi.org/10.13039/501100004281) grant Sonata to ML no. 2018/31/D/ST3/02408.

### v3

@@ I Introduction
In many applications, the dynamics of the hard degrees of freedom are of secondary importance, and one method of circumventing this difficulty is to treat them as fully constrained (i.e. rigid), thus eliminating the short timescale, and enabling faster simulations. This is the basis of many algorithms, such as SHAKE [4] and its further extensions. During the development of such rigid models, a disconnect was discovered between the equilibrium bond angle distributions in the rigid simulations and the simulations with bonds modelled as very stiff springs. Further summary of literature on the topic is discussed in Sec. II.
@@
In this contribution, we present a clear mathematical procedure leading to the $\mathrm{d}p_{k\to\infty}$ limit in the overdamped regime when inertial effects can be neglected. We show that the distribution given by Eq. (1) is not completely determined by fixing the bond lengths. To correctly predict the limiting distribution, knowledge about the nature of confinement is needed, and its structure survives in the final expression. Furthermore, we demonstrate that Fixman’s expressions for the stiff limit miss a term, which is important even in the case of harmonic springs.
@@
Fixman [5] presented a purely classical calculation of the limiting distribution, correctly writing the action of the distribution on a test function but erroneously assuming that the constraining potential depends only on the hard coordinates. As a result, the details of the confining potential cannot appear in his final expressions, which contain only the projected phase-space volume element terms. Gō and Sheraga [6] treated the problem quantum-mechanically and correctly identified two contributions to the equilibrium probability density of the soft coordinates: the projected volume elements and the zero-point energy of the vibrational motion of the hard coordinates. However, they then stated that the zero-point-energy contribution in typical molecular systems varies only a little when changing the soft coordinates and may be altogether neglected. We show in the following that this is not necessarily the case in overdamped dynamics. By focusing on a concrete classical two-dimensional example with one constraint, Helfand [11] correctly identified the rigid-rod type distribution as a uniform distribution on the constraining manifold and computed the stiff-spring limit in a similar simple case. Unfortunately, these results are not expressed in a form applicable to higher-dimensional examples with multiple constraints.
@@
Hinch et al. [8, 12] approached the problem differently, by manipulating the Langevin equations, which are notoriously difficult to handle, and erroneously assumed that details of the confining potential are of little importance to the equilibrium distribution of the constrained configurations. Van Kampen and Lodder [13] noted that the approach of Helfand [11] is generally applicable to constrained systems and comment that the discrepancy between the stiff-spring and rigid-rod distributions in the case of the trimer molecule is due to the fact that the width of this gully is not the same everywhere [13]. However, they did not quantify the effect of this width in a higher-dimensional case. As a result, the knowledge of the influence of the shape of the constraining potential appears to have been lost as multiple works (for example, [14, 15, 3]) mention only the densities of the volume elements and omit the zero point energy or, equivalently, the potential shape terms from their descriptions of constrained dynamics, widely using the result of Fixman [5] as a reference.
In the following, we show a way to compute the limiting distribution by integrating out the confined degrees of freedom in the overdamped limit. We explicitly find the metric contribution arising from the transformation of coordinates and the shape term that represents the details of the constraining potential and which remains imprinted on the configurational distribution when the hard coordinates are taken to be infinitely stiff.
@@ III The stiff spring limit
This work focusses on the distributional limit of equilibrium distributions at increasingly stiff confinement: the stiff spring limit. We begin our consideration with a careful definition of this limit. The rigid rod distribution is discussed only insofar as it is needed for the “ratio of densities” argument presented by Frenkel and Smit [3].
We present a derivation of the limiting distribution in the following setting, sketched in Fig. 1. Consider a molecule (or a polymer) with a handful of identical subunits (atoms, beads, or monomers) each of mass $m$ and a total of $2N$ degrees of freedom ($N$ configurational and $N$ momentum) in a heat bath at constant temperature $T$, with $\beta=1/k_{B}T$. The conformation of the molecule is described by an $N$-dimensional vector $\bm{q}$ and the velocities of monomers are specified by a vector $\bm{v}$. We parametrise the conformations by Cartesian coordinates of monomers $q_{i}$ and Cartesian components of velocities $v_{i}$ with $i\in 1\dots N$. The potential energy of the molecule has two components: the conformation-dependent energy $U(\bm{q})$ and the confining (springs) potential $k^{2}W(\bm{q})$, where $k$ is a large parameter describing the spring stiffness. Note that $k$ here would correspond to the square root of a harmonic spring constant. The probability density $\hat{p}_{k}$ of the equilibrium distribution of the system configuration with respect to the Lebesgue measure $\mathrm{d}\bm{q}\mathrm{d}\bm{v}$ is given by the Maxwell-Boltzmann distribution [16]
@@
with a (different) normalising function $N(k)$. Crucially, this density is given with respect to canonical conformation coordinates and not just any coordinates – if generalised coordinates are used, more care is needed and volume element density terms appear in expressions analogous to Eq. (4). Moreover, in the case of generalised coordinates, the expression for kinetic energy is also more complicated. Clearly, Eq. (4) is a physical object and we are free to work in any coordinates of choice, and for the purpose of the derivation of the stiff spring limit Cartesian coordinates are particularly convenient. We return to the treatment of generalized coordinates in later in this paper.
@@
$I_{k}=\iint\frac{N(k)}{k^{N-M}}\exp\left\{-\beta U(\bm{r},k^{-1}\bm{t})-\beta\left[\bm{t}^{T}\bm{\mathsf{H}}(\bm{r})\bm{t}+||\bm{t}||^{2}\frac{o(k^{-2}||\bm{t}||^{2})}{k^{-2}||\bm{t}||^{2}}\right]\right\}\phi(\bm{r},k^{-1}\bm{t})\mathrm{d}\bm{t}\mathrm{d}\bm{r}.$ (9)
@@
$I=\iint\left(\lim_{k\to\infty}\frac{N(k)}{k^{N-M}}\right)\exp\left[-\beta U(\bm{r},\bm{0})-\beta\,\bm{t}^{T}\bm{\mathsf{H}}(\bm{r})\bm{t}\right]\phi(\bm{r},\bm{0})\mathrm{d}\bm{t}\mathrm{d}\bm{r},$ (10)
@@
$L=\lim_{k\to\infty}\frac{N(k)\sqrt{(2\pi\beta)^{N-M}}}{k^{N-M}},$ (12)
@@
for an appropriate constant $L$, and since $\phi$ is arbitrary, we arrive at $p_{\infty}=\lim_{k\to\infty}p_{k}$ such that
@@
in the weak sense. We have thus arrived at the central result of this paper.
@@
IV Calculation of the shape term
@@
V Calculation for the trimer
Figure 2: (a) Parametrisation of the trimer uses the position of the central particle ($b$), bond angle $\psi$ and bond lengths $l_{a}$ and $l_{b}$, with identical equilibrium spring lengths $l_{0}$. Rigid-body rotation around the particle $b$ is omitted for clarity. (b) Motion along the symmetric (top) and antisymmetric eigenvectors of the hessian $\bm{\mathsf{H}}$ of the harmonic confinement function. Note that the central bead must move for the overall motion to be orthogonal to the constraining manifold $\mathcal{M}$. (c) Equilibrium probability density function (PDF) for the bond angle $\psi$ from the Langevin numerical simulation. Solid line is the theoretical prediction of Eq. (34), dashed line is the uniform distibution given by projection of canonical measure given by Eq. (29).
@@ V Calculation for the trimer
from the matrix $\bm{\mathsf{J}}$ listed in the Appendix. Next, we compute the shape term using Eq. (19). Intermediate steps of the calculation require computing a 2-by-9 matrix $\bm{\mathsf{A}}$ which has a number of zero entries but contains lengthy trigonometric expressions. These simplify somewhat for points on $\mathcal{M}$, but were handled with symbolic computation engine Mathematica. Computing 2-by-2 matrices $\bm{\mathsf{A}}^{T}\bm{\mathsf{A}}$ and $\bm{\mathsf{A}}^{T}\bm{\mathsf{B}}\bm{\mathsf{A}}$, after simplification, gives compact closed-form expressions, which, as expected, depend only on the angle $\psi$, via
@@
where $N$ is a normalisation constant. Therefore, the marginal density of the bond angle is proportional to $\sin\psi$ as expected. However, this is simply not the case of a uniform distribution––that distribution corresponds to $|\bm{\mathsf{H}}|=\text{const}$ and the notion of uniformity is inherited from the canonical measure in the canonical (Cartesian) coordinates displayed as uniform distribution in Fig. 2(c). This can be examined in greater detail by looking at the eigenvectors of $\bm{\mathsf{H}}$, visualised in Fig. 2(b). There are two eigenvectors which correspond to symmetric and antisymmetric motion of the terminal beads. Counterintuitively, the vectors in $\mathcal{M}^{\perp}$ change the orientation of at least one bond when the bond length changes.
@@
VI Cyclic tetramer
Figure 3: (a) The shape of a cyclic tetramer is specified with the bead $b$ taken as reference point. The directions of sections $ab$ and $bc$ span a rhombus, which is used to define the position of bead $d$. (b) Equilibrium probability density function (PDF) for the bond angle $\psi$ from the numerical simulation. The solid line is the theoretical prediction of Eq. (49). The dashed line is the uniform distribution obtained by projecting the canonical (Cartesian) measure.
@@ VI Cyclic tetramer
VII Conclusions
In this work, we have provided a procedure by which one can find the equilibrium probability probability distribution function of the configurations of a bead-spring model of a polymer in the overdamped regime in the limit of infinite stiffness of the bonds. In particular, we have demonstrated that the shape of the confining spring potential persists in the limiting distribution. By rephrasing the problem as a mathematically rigorous limit, we have shown that the classical expression given by Fixman [5], later reproduced in numerous works and books, for example in Ref. [3], is missing a critical term describing the shape of the confining potential, $|\bm{\mathsf{H}}|$. The presented reasoning can be applied without changes to potentials which depend in an eventually monotone way on the large parameter $k$, dropping the assumption of linear dependence on $k$. Furthermore, in the case of a very flat and very sharp confining potentials, which do not admit a second-order Taylor expansion with an almost-surely non-vanishing Hessian, one can sometimes reduce the problem to the presented case by a change variables in the soft coordinates such as $s\to s^{\gamma}$. For other, badly-behaved confining potentials, such as $1/\log(1/s)$, a different methodology is required.
Our work shows that the reasoning presented by Frenkel & Smit [3] cannot be universally applied. First, the dynamics and the distribution of a polymer connected with (stiff or soft) springs is independent of the mass of the monomers in the overdamped regime, and thus the limiting distribution cannot depend on the mass of the monomers. Second, putting aside the question of existence of the rigid rod distribution (or the lack thereof), the supposition that the stiff spring limit should always be the same, regardless of the details of the confining potential is clearly not true, and harmonic potentials will lead to a different outcome than springs that realise uniform confinement around the constraining manifold, as previously remarked, e.g., by van Kampen & Lodder [13], although without a general mathematical formulation. Thus it cannot be true that the ratio of densities of the rigid rod distribution to the stiff spring limit distribution is given by the manifold $\mathcal{M}$ alone. The difference between uniform confinement and harmonic spring confinement can be arbitrarily large, as shown in the tetramer example. Frenkel & Smit remark in their book that for the bond length constraints of the type most often used in MD simulations, the effect of hard constraints on the distribution functions seems to be relatively small [3, Chapter 15.1.1]. We believe this statement to be misleading. In the discussed case of the trimer molecule, the effect might indeed be small, but we have shown that for the tetramer system, which is still a relatively simple setup, the effect is pronounced. To provide practical insight, we propose a technique for the analysis of limiting distributions once the confining potential is specified in detail.
@@ Acknowledgements.
The Authors thank Bartłomiej Lewandowski for his insightful comments regarding convergence of distributions. The work of ML and RW was supported by the National Science Centre of Poland (FundRef DOI: http://dx.doi.org/10.13039/501100004281) grant Sonata to ML no. 2018/31/D/ST3/02408.